OpenStudy (anonymous):

you have 4/6 cup of flour in your cupboard.A recipe for bread calls for 1/4 cup of flour.How nuch flour would you have left it you made the bread

OpenStudy (anonymous):

1/4 * 6/6 = 6/24 4/6* 4/4 = 16/24 16/24 - 6/24 = 10/24 which can be simplified down to 5/12

OpenStudy (anonymous):

a recipe for nut bbread calls 1/3 cup of walnuts and 2/6 cup of pecans.what is the total amount of wallnuts and pecans needed for the recipe

OpenStudy (anonymous):

walnuts : 1/3 * 2/2 = 2/6 pecans: 2/6 2/6 + 2/6 = 4/6, simplified would be 2/3 (:

OpenStudy (anonymous):

thank you ...ben is making waffles for his 3 friends.the recipe calls for 2/3 cup of bleached flour and 1/4 cup of wheat flour.what is the total ammount of flour ben needs to make the waffles?

OpenStudy (anonymous):

what operation is needed to sove that problem

OpenStudy (anonymous):

bleached: 2/3 * 4/4 = 8/12 wheat: 1/4 * 3/3 = 3/12 8/12+3/12 = 11/12 cups

OpenStudy (anonymous):

what are the options?

OpenStudy (anonymous):

A./ B.+ C.- D.*

OpenStudy (anonymous):

that question is weird /: you need to multiply the fractions to get a common denominator before you add them, and then you need to add them. Im guessing its multiplication because addition is obvious. but its up to you ^.^

OpenStudy (anonymous):

it was +, but ben previoiusly made waffles for his 3 friends.the recipe called for 2/3 cup of bleached flour and 1/4 cup of wheat flour.if ben is making waffles for 5 friends,he might need to ______the recipe and, how old are you,not to be noisy just cause you are so smart

OpenStudy (anonymous):

oops sorry hehe. what are the choices for this one? it sounds like it should be double, increase, multiply. Im 16 almost 17 lol thank you (:

OpenStudy (anonymous):

i had guessed and it was double but ,you have 3/6 cup of oatmeal in your cupboard.Arecipe for cookies calls for 1/4 cup of oatmeal would you have left if you made the cookies

OpenStudy (anonymous):

have: 3/6 * 4/4 = 12/24 need: 1/4 * 6/6 = 6/24 12/24 - 6/24 = 6/24 which simplifies to 1/4 (:

OpenStudy (anonymous):

thanks,you have 3/6 cup of cornmeal in your cupboard.A recipe for muffins calls for 1/4 cup of coernmeal.how much cornmeal would you have left if you made the muffins?

OpenStudy (anonymous):

same as the above question haha

OpenStudy (anonymous):

ohh silly meh,A recipe for a pie calls for 1/3 cup of blackberries and 2/6 cup of rasberries.What is the total amount of berries needed for the recipe?

OpenStudy (anonymous):

black: 1/3 * 2/2 = 2/6 ras: 2/6 4/6 or 2/3

OpenStudy (anonymous):

ben is making waffles for his 3 friends.the recipe calls for 2/3 cup of bleached flour and 1/4 cup of wheat flour.what is the total amount of flour ben needs to make waffles